Solve for x : cot(sinx+3)=1

Grade:12th Pass

2 Answers

Ameena Manzoor
9 years ago
cot(sinx+3)=1
cot(sinx+3)=cot(pi/2)
Taking reciprocal,
tan(sinx+3)=tan(pi/2)
This is of the form tanx=tany,
Hence we apply our formula for gen. solution,
Therefore, sinx+3 = n(pi) + (pi)/4
Hence x = sin inverse (5(pi)/4 – 3)
 
Yash Baheti IIT Roorkee
9 years ago
Hi,

Cot-11= pi/4
Sin x + 3 = pi/4
Sin x = pi/4 – 3
| pi/4 – 3 | > 1
So no solution